# Deep Reinforcement Learning for Time-Critical Wilderness Search And Rescue Using Drones

### v1

@@ 1 Introduction
Another class of algorithm that has surpassed human abilities in highly complex tasks is Deep Reinforcement Learning (DRL), a subset of machine learning. DRL has been successfully applied to playing video games Mnih et al. [2015], OpenAI [2019] and achieving championship-level drone racing Kaufmann et al. [2023]. The ability of DRL to generalize the problem and make decisions based on extensive training allows it to provide unique solutions that other algorithms may not be able to achieve.
At the core of the problem, WiSAR mission planning involves the effective use of any a priori data, such as the place last seen, age, and fitness levels of the missing person. This information can be used to generate a Probability Distribution Map (PDM) Ewers et al. [2023a], Hashimoto et al. [2022], Šerić et al. [2021], which describes the probability of detecting the missing person at a given location and informs the search stage of the mission.
@@
Further related work is discussed in Section 2, and the methodology is presented in Section 3. Results are shown in Section 4 with a conclusion being drawn in Section 6.
@@ 3 Method
3.1 The Agent
The agent is modelled as a simple heading control model with a constant step size $s$. It is assumed that any vehicle executing this mission can perfectly track the waypoints through its controller or operator. Thus, the position vector $\mathbf{x}\in\mathbb{R}^{2}$ is updated via
$\mathbf{x}_{t+1}=\mathbf{x}_{t}+s\begin{bmatrix}\cos u_{t}\\ \sin u_{t}\end{bmatrix}$ (1)
@@ 3.2 PDM Modelling
where $\mathbf{\mu}_{i}$ and $\mathbf{\sigma}_{i}$ are the mean location and covariance matrix of the $i$th bivariate Gaussian respectively. If the bounding area were infinite, that is $x_{\text{min}}=y_{\text{min}}=-\infty$\mathrm{m}$$ and $x_{\text{max}}=y_{\text{max}}=\infty$\mathrm{m}$$, then $\sum p(\mathbf{x})=1$. However, as can be seen from Figure 2, the area enclosed by the rectangular bounds contain less than this. Section 3.3 further discusses how this is handled such that the available probability is normalized.
@@ 3.3 Reward Modelling
$k$ $=\frac{a_{\textit{area}}}{2(\frac{1}{2}\pi{R_{\textit{buffer}}}^{2})+2{R_{\textit{buffer}}}s}$ (8)
$=\frac{a_{\textit{area}}}{{R_{\textit{buffer}}}(\pi{R_{\textit{buffer}}}+2s)}$ (9)
@@
Figure 4: The anatomy of the area calculation of an isolated step of length $s$ with buffer ${R_{\textit{buffer}}}$ as used in Equation 9.
As highlighted in Section 3.2, the enclosed probability by the bounds is not equal to $1$. To handle this, $\Delta p_{t}$ is scaled by the total available probability within the search area $p_{A}=I(A)$. Combining $p_{A}$ with Equation 7 and Equation 9, gives the reward
@@ 3.4 Training Algorithm
SACHaarnoja et al. [2018] is a DRL algorithm well-suited for continuous control problems like drone navigationKaufmann et al. [2023]. It tackles the challenge of exploration vs exploitation by simultaneously maximizing expected reward and entropy. Entropy encourages exploration of the environment, preventing the policy from getting stuck in local optima. SAC achieves this by learning a policy, multiple Q-functions to evaluate actions, and a temperature parameterHaarnoja et al. [2019] that controls the trade-off between exploration and exploitation.
@@ 3.4.1 Policy Architecture
The policy architecture designed to handle this observation is defined in Figure 5. The sub-FE for the $s_{\text{path}}$ state observation is of particular note as it uses a 2D CNN architecture based on work by Mnih et al. [2015].
@@ 4.1 Experimental setup
In order to effectively benchmark the proposed algorithms, two additional baselines are implemented; lawnmowerGalceran and Carreras [2013] and LHC_GW_CONV Lin and Goodrich [2009]. These were chosen due to the former being ubiquitous for coverage planning, and the latter being a optimisation-based implementation that struggles to fully explore the PDM. To ensure compatibility in the comparison to the proposed algorithm, the parallel lines for lawnmower are offset by the step size $s$ and the grid dimensions for LHC_GW_CONV  are $(x_{\text{max}}-x_{\text{min}},y_{\text{max}}-y_{\text{min}})/s$. The maximum number of waypoints $N_{\text{waypoint}}$ are converted to a maximum distance $D_{\text{max}}=sN_{\text{waypoint}}$ and the generated paths are truncated at this point.
@@
One evaluation of an algorithm involves generating the random PDM, then creating the resultant search path. This is labelled one run. Table 2 shows the number of runs undertaken per algorithm and is the base of the following analysis.

### v2

@@ 1 Introduction
Another class of algorithm that has surpassed human abilities in highly complex tasks is Deep Reinforcement Learning (DRL), a subset of machine learning. DRL has been successfully applied to playing video games Mnih et al. [2015], OpenAI [2019] and has achieved championship-level drone racing Kaufmann et al. [2023]. The ability of DRL to generalize the problem and make decisions based on extensive training allows it to provide unique solutions that other algorithms may not be able to achieve.
At the core of the problem, WiSAR mission planning involves the effective use of a priori data, such as the place last seen, age, and fitness levels of the missing person. This information can be used to generate a Probability Distribution Map (PDM) Ewers et al. [2023a], Hashimoto et al. [2022], Šerić et al. [2021], which describes the probability of detecting the missing person at a given location and informs the search stage of the mission.
@@
Further related work is discussed in Section 2, and the methodology is presented in Section 3. Results are shown in Section 4 with limitations being discussed in Section 5, and a conclusion is drawn in Section 6.
@@ 3 Method
3.1 Modelling
3.1.1 Environment
The drone within the environment is modelled as a simple heading control model with a constant step size $\lambda$. It is assumed that any drone executing this mission can perfectly track the waypoints through its controller or operator. Thus, the position vector $\mathbf{x}\in\mathbb{R}^{2}$ is updated via
$\mathbf{x}_{t+1}=\mathbf{x}_{t}+\lambda\begin{bmatrix}\cos u_{t}\\ \sin u_{t}\end{bmatrix}$ (1)
@@ 3.1.2 PDM
where $\mathbf{\mu}_{i}$ and $\mathbf{\sigma}_{i}$ are the mean location and covariance matrix of the $i$th bivariate Gaussian respectively. If the bounding area were infinite, that is $x_{\text{min}}=y_{\text{min}}=-\infty$\mathrm{m}$$ and $x_{\text{max}}=y_{\text{max}}=\infty$\mathrm{m}$$, then $\sum p(\mathbf{x})=1$. However, as can be seen from Figure 1, the area enclosed by the rectangular bounds contain less than this. Section 3.1.3 further discusses how this is handled such that the available probability is normalized.
@@ 3.1.3 Reward
$k$ $=\frac{a_{\textit{area}}}{2(\frac{1}{2}\pi{R_{\textit{buffer}}}^{2})+2{R_{\textit{buffer}}}\lambda}$ (8)
$=\frac{a_{\textit{area}}}{{R_{\textit{buffer}}}(\pi{R_{\textit{buffer}}}+2\lambda)}$ (9)
@@
Figure 3: The anatomy of the area calculation of an isolated step of length $\lambda$ with buffer ${R_{\textit{buffer}}}$ as used in Equation 9.
As highlighted in Section 3.1.2, the enclosed probability by the bounds is not equal to $1$. To handle this, $\Delta p_{t}$ is scaled by the total available probability within the search area $p_{A}=I(A)$. Combining $p_{A}$ with Equation 7 and Equation 9, gives the reward
@@ 3.2 Training Algorithm
SACHaarnoja et al. [2018] is a DRL algorithm well-suited for continuous control problems like drone navigationKaufmann et al. [2023]. It tackles the challenge of exploration vs exploitation by simultaneously maximizing expected reward and entropy. Entropy encourages exploration of the environment, preventing the policy from getting stuck in local optima. SAC achieves this by learning a policy, multiple Q-functions to evaluate actions, and a temperature parameterHaarnoja et al. [2019] that controls the trade-off between exploration and exploitation. Figure 4 outlines the iterative process of reinforcement learning, where the policy is used to generate actions $a_{t}$ and the environment then provides the resultant observation $s_{t}$ and reward $r_{t}$.
@@ 3.2.1 Policy Architecture
The policy architecture designed to handle this observation is defined in Figure 5. The sub-FE for the $r_{\text{path}}$ state observation is of particular note as it uses a 2D CNN architecture based on work by Mnih et al. [2015].
@@ 4.1 Experimental setup
In order to effectively benchmark the proposed algorithms, two additional baselines are implemented; lawnmowerGalceran and Carreras [2013] and LHC_GW_CONV Lin and Goodrich [2009]. These were chosen due to the former being ubiquitous for coverage planning, and the latter being a optimisation-based implementation that struggles to fully explore the PDM. To ensure compatibility in the comparison to the proposed algorithm, the parallel lines for lawnmower are offset by the step size $\lambda$ and the grid dimensions for LHC_GW_CONV  are $(x_{\text{max}}-x_{\text{min}},y_{\text{max}}-y_{\text{min}})/\lambda$. The maximum number of waypoints $N_{\text{waypoint}}$ are converted to a maximum distance $D_{\text{max}}=\lambda N_{\text{waypoint}}$ and the generated paths are truncated at this point.
@@
One evaluation of an algorithm involves generating the random PDM, then creating the resultant search path. This is labelled one run. Table 2 shows the number of runs undertaken per algorithm and this generated data is base of the following analysis.